1. A kettle uses 167200J of heat energy when it heats water from 200C to 1000C.  How much water is in the kettle? The specific heat capacity of water is 4180j/kg0C.


  1. The temperature of a 0.8 kg metal block is raised from 27 °C to 77 °C when 4200 J of energy is supplied.  Find the specific heat capacity of the metal.


  1. 9600 J of heat energy is supplied to 1 kg of methylated spirit in a polystyrene cup.


Calculate the rise in temperature produced.

Take the specific heat capacity of methylated spirit to be 2300 J/kg°C.


  1. A 750g steel cooking pan cools from 650C to 180C.  How much heat energy is released by the pan?  The specific heat capacity of steel is 500j/kg0C.


  1. An insulated container contains 0.5kg of water at an initial temperature of 160C.  A heater is placed in the water and switched on.  It supplies 83600J to the water.  Calculate the final temperature of the water.  (Specific heat capacity of water = 4180J/kg0C)




1
Expert's answer
2021-05-12T08:38:20-0400

1.

"m=\frac{Q}{c(t_2-t_1)}=0.5~kg."

2.

"c=\frac{Q}{m(t_2-t_1)}=105~\frac{J}{kg\cdot °C}."

3.

"\Delta t=\frac{Q}{mc}=4°C."

4.

"Q=cm(t_2-t_1)=17625~J."

5.

"t_2=t_1+\frac{Q}{cm}=56°C."